The concentration of NaCl in MSA is 7.5%. a) What is the prupose of this high NaCl concentration? Give a detailed explanation. b) What is the molarity of NaCl in this medium? Show your calculation. (M.W. of NaCl 58.44)
Answer #1

Answer:

Mannitol salt agar contains a high concentration NaCl (7.5%). This makes it selective media for Gram-positive bacteria (Staphylococcus and Micrococcaceae).

It inhibits the growth of other bacteria.

Molarity of NaCl

7.5g per 100mL

75g in 1000mL

Number of moles= weight in g/Mol wt = 7.5/58.44 =1.2

Molarity= number of moles/volume in litre = 1.2/1L = 1.2 M